Does the equation 2/r= 3cos(θ-π/4)+ 2sin(θ+π/4) represent a straight line? Satisfy your answer
1
Expert's answer
2020-11-19T17:01:58-0500

"\displaystyle


\frac{2}{r}= 3\cos\left(\theta -\frac{\pi}{4}\right)
+ 2\sin\left(\theta +\frac{\pi}{4}\right)\\

r = \frac{2}{3\cos\left(\theta -\frac{\pi}{4}\right)
+ 2\sin\left(\theta +\frac{\pi}{4}\right)}\\

\cos\left(\frac{\pi}{2} - \theta\right) = \sin{\theta}\\

r = \frac{2}{3\sin\left(-\theta + \frac{3\pi}{4}\right)
+ 2\sin\left(\theta +\frac{\pi}{4}\right)}\\


\sin{\theta} = \sin(\pi - \theta)\\

r = \frac{2}{3\sin\left(\theta + \frac{\pi}{4}\right)
+ 2\sin\left(\theta +\frac{\pi}{4}\right)}\\



r = \frac{2}{5\sin\left(\theta + \frac{\pi}{4}\right)}\\


r^2 = \frac{4}{25\sin^2\left(\theta + \frac{\pi}{4}\right)}\\


x^2 + y^2 = \frac{4}{25\sin^2\left(\theta + \frac{\pi}{4}\right)}\\


x^2 + y^2 = \frac{8}{25\left(1 - \cos\left(2\theta + \frac{\pi}{2}\right)\right)}\\


x^2 + y^2 = \frac{8}{25\left(1 + \sin(2\theta)\right)}\\


x^2 + y^2 = \frac{8}{25\left(1 + 2\sin{\theta}\cos{\theta}\right)}\\

x = r\cos\theta, y = r\sin\theta\\

x^2 + y^2 = \frac{8}{25\left(1 + \frac{2xy}{r^2}
\right)}\\


x^2 + y^2 = \frac{8(x^2 + y^2)}{25\left(x^2 + y^2 + 2xy
\right)}\\



1 = \frac{8}{25\left(x^2 + y^2 + 2xy
\right)}\\


x^2 + y^2 + 2xy = \frac{8}{25}\\


(x + y)^2 = \frac{8}{25}\\


x + y = \pm\frac{2\sqrt{2}}{5}\\


\textsf{This is of course the}\\
\textsf{equation of a straight line.}\\




\textsf{A plot is attached above}"